# UltraViCo: Breaking Extrapolation Limits in Video Diffusion Transformers

### v1

@@ 1 Introduction
Building upon the expressive power of diffusion transformers (DiTs) (bao2023all; peebles2023scalable), recent advances in text-to-video (T2V) generation bao2024vidu; zheng2024opensora; videoworldsimulators2024; wan2025wan; kong2024hunyuanvideo; hong2022cogvideo have enabled models to synthesize high-fidelity videos. However, these models are typically trained on a fixed maximum sequence length (e.g., 5 seconds wan2025wan; kong2024hunyuanvideo; hong2022cogvideo) and struggle to generate videos beyond their training length, a task we term video length extrapolation, which is critical for practical applications.
To investigate the core challenges of this task, we conduct experiments on a range of models and identify two failure modes: (i) a model-specific periodic content repetition, where short clips loop indefinitely in certain models; and (ii) a universal quality degradation, manifested as blurred spatial details and frozen temporal dynamics across all models. Both failures become increasingly severe as the extrapolation length grows. Prior work, such as RIFLEx (zhao2025riflex), tackles repetition from the perspective of positional encodings, while overlooking quality degradation and therefore achieving limited extrapolation. We contend, however, that positional encodings play only an indirect role by perturbing queries and keys to influence attention. In contrast, attention itself—directly aggregating contextual information to generate outputs—offers a more fundamental view.
@@
To validate our approach, we conduct comprehensive evaluations on various T2V models (kong2024hunyuanvideo; yang2024cogvideox; wan2025wan) and extrapolation ratios, against a large family of baselines (chen2023extending; bloc97; zhuo2024lumina; peng2023yarn; zhao2025riflex). Experiments demonstrate that our method consistently surpasses all baselines in all settings by simultaneously addressing both failure modes. Notably, while prior methods collapse beyond $3\times$ extrapolation and yield static videos, ours maintains fluid motion, effectively extending the practical limit from $2\times$ to $4\times$. Remarkably, it improves Dynamic Degree and Imaging Quality by 233% and 40.5% over the previous best method at $4\times$ extrapolation. Beyond this, our method also generalizes seamlessly to downstream tasks such as various controllable video synthesis and editing.
HunyuanVideo
Wan
Normal
length
Refer to caption Refer to caption
@@
3 $\times$
extra.
Refer to caption Refer to caption
@@ 3 $\times$
Variable
extra.
@@ 2 Preliminary
Attention mechanism with rotary position embedding. Modern video diffusion models are largely built on DiTs whose core is the attention mechanism (vaswani2017attention). The input video is patched into $L$ tokens, each projected into queries, keys, and values. To encode the position information, DiTs mainly adopt Rotary Position Embedding (RoPE) (su2024roformer), which injects position into queries and keys through complex rotations. Concretely, for each query or key vector ${\bm{x}}\in\mathbb{R}^{D}$ at position $t$, RoPE maps it to $\mathbb{R}^{D}$ as
@@
For videos with temporal and spatial axes, Multimodal RoPE (M-RoPE) (wang2024qwen2) partitions the dimension $D=d_{\mathcal{T}}+d_{\mathcal{H}}+d_{\mathcal{W}}$ and encodes each subspace separately. Since we focus on temporal extrapolation, we consider only the temporal axis and denote $d_{\mathcal{T}}$ as $d$ for simplicity (see details in Appendix B.2).
Problem setting: video length extrapolation. Despite advances, DiT-based video generation models struggle to produce videos longer than their training duration. This task, known as video length extrapolation  (zhao2025riflex), aims to adapt a pre-trained model to generate high-quality videos of a sequence length $L^{\prime}$ that exceeds its training length $L$, with the extrapolation ratio defined as $s=L^{\prime}/L>1$. Notably, video length extrapolation targets the model’s intrinsic ability to generate longer sequences in a single forward generation, which is orthogonal to prior methods (qiu2023freenoise; wang2023gen; kim2024fifo; wang2024loong; lu2024freelong) that rely on inference-time modifications. See Appendix A for more related work.
@@ 3.1 Failure Modes of Video Length Extrapolation
In this section, we investigate the core challenges of video length extrapolation on a range of SOTA video diffusion transformers, including Wan (wan2025wan), HunyuanVideo (kong2024hunyuanvideo), and CogVideoX (yang2024cogvideox).
Qualitative results in Fig.2a and Fig.2b reveal two distinct failure modes. The first is a periodic content repetition, which occurs in certain models such as HunyuanVideo and CogVideoX. The second is a universal quality degradation, characterized by compromised spatial fidelity and temporal dynamics across all models. To further investigate their trends across extrapolation lengths, we perform a quantitative analysis on 10 prompts using metrics including Imaging Quality (huang2024vbench), Dynamic Degree (huang2024vbench), and Repetition Count. Fig. 2c confirms that both failures become more severe as the extrapolation factor increases.
@@ 3.2.1 The Cause of Content Repetition: Periodic Attention Patterns
Model
Attention maps
Statistical row-wise attention analysis
Hun.
Refer to caption Refer to caption
@@
Wan
Refer to caption Refer to caption
@@
Model
Generated videos: baseline vs. intervention
Attention maps: baseline vs. intervention
Hun.
Refer to caption Refer to caption
@@
Wan
Refer to caption Refer to caption
@@ 3.3 UltraViCo
Efficient CUDA implementation. UltraViCo requires modifying attention logits, but standard PyTorch attention is infeasible for long sequences. At a $3\times$ extrapolation ($\sim$200K tokens for HunyuanVideo), for instance, materializing a $200\text{K}\times 200\text{K}$ attention mask consumes over 8080GB of memory in bf16, causing an immediate out-of-memory error. To address this, we integrate UltraViCo into Triton-based FlashAttention (dao2022flashattention) and SageAttention (zhang2024sageattention), where the online-softmax formulation avoids explicit mask construction. This yields scalable, memory-efficient computation, enabling UltraViCo on large video models.
@@ 4.1 Setup
We evaluate methods on three video diffusion models, including HunyuanVideo, Wan2.1-1.3B and CogVideoX-5B. Following RIFLEx, we use 100 prompts sampled from VBench (huang2024vbench). For quantitative evaluation, following RIFLEx, we adopt Imaging Quality (Quality), Dynamic Degree (Dynamics), and Overall Consistency (Overall) from VBench, along with the NoRepeat Score for models prone to content repetition. Notably, our NoRepeat Score is a variant of that in RIFLEx, tailored for multiple-repetition (see Appendix C.2 for details). Finally, we conduct a user study with 10 participants on 10 prompts, where users rank (User) the overall quality of videos across all methods. More details are provided in Appendix C.2.
@@ 4.2 Results
Performance comparison. We compare a wide range of length extrapolation baselines on three SOTA models (kong2024hunyuanvideo; yang2024cogvideox; wan2025wan) across various extrapolation ratios, including PE (zhao2025riflex), PI (chen2023extending), NTK (bloc97), TASR (zhuo2024lumina), YaRN (peng2023yarn), and RIFLEx. Tab. 1 reports $3\times$ and $4\times$ results on HunyuanVideo and Wan, while Fig. 6 shows qualitative samples on HunyuanVideo. Results for additional ratios and models are provided in the Appendix C.3.
@@
Connection with other long-video generation methods. UltraViCo aims to extend the effective training window of video diffusion transformers and is therefore orthogonal to existing long-video generation techniques such as FreeNoise (qiu2023freenoise), FIFO-Diffusion (kim2024fifo), and sliding-window. As demonstrated in Table 2, enlarging the context window via UltraViCo consistently improves the long-term temporal consistency of these methods, without negatively affecting other performance. In Table 2, all methods follow the same evaluation setup ($6\times$ extrapolation for 30-second videos on Wan), where UltraViCo extends the base model’s training window by $3\times$.
Generalization to downstream tasks. Our method enhances the model’s inherent ability to handle longer sequences, making it naturally applicable to downstream tasks. As shown in Fig. 1, based on VACE (jiang2025vace), UltraViCo enables $3\times$ extrapolation in controllable generation and video editing. See Appendix C.4 for additional results.
@@ 5 Conclusion
Reproducibility STATEMENT
Our code and the prompts in the paper are included in the supplementary material, and the implementation details are described in Sec. 4.1.
Use of Large Language Models
We used a large language model solely to assist in polishing English writing and improving clarity. All research ideas, experiments, results, and interpretations are entirely our own.
@@ Appendix A Related Work
The recent advances in text-to-video generation have been primarily driven by diffusion models (ho2020denoising; song2020score; ho2022imagen; he2022lvdm; zhao2022egsde; zhao2023controlvideo; blattmann2023stable; xing2023dynamicrafter; chen2023videocrafter1; zhao2024identifying; polyak2024movie; zhou2024allegro; genmo2024mochi; chen2024videocrafter2). With the development of diffusion transformers (DiTs) (bao2023all; peebles2023scalable), DiT-based text-to-video diffusion models have achieved remarkable performance, such as Sora (videoworldsimulators2024), Vidu (bao2024vidu), CogVideoX (yang2024cogvideox) and Open-Sora (zheng2024open). Although achieving high quality, leading models are trained only on a fixed maximum sequence length, limiting long-term capacity. During video length extrapolation, they suffer from repetition or quality degradation, underscoring the need for length extrapolation.
@@
The goal of length extrapolation is to enable transformers to generate sequences longer than those seen during training in a single forward (press2021train). This is typically achieved by modifying positional encodings. For example, position interpolation (PI) (chen2023extending) improves performance by interpolating the frequencies in RoPE so that they remain within the training range even under extrapolation. NTK (bloc97), YaRN (peng2023yarn), and Time-aware Scaled RoPE (TASR) (zhuo2024lumina) combine interpolation with direct extrapolation, incorporating adjustments along the token dimension, denoising timesteps, and other factors to achieve better results. However, these methods perform poorly on image and video DiTs, often leading to content collapse or repetition. RIFLEx (zhao2025riflex) mitigates repetition by identifying and attenuating the intrinsic RoPE frequency, yet it still suffers from degraded visual quality. In contrast, our method effectively addresses both content repetition and quality degradation.
@@
There also exist many approaches to long video generation (qiu2023freenoise; wang2023gen; henschel2025streamingt2v; kim2024fifo; tan2024video; yin2025slow; wang2024loong; cai2025ditctrl; li2025longdiff; lu2024freelong; tan2025freepca; jiang2025lovic; gao2025longvie; gu2025long), most of which intervene in the diffusion inference process. For instance, FreeNoise (qiu2023freenoise) enhances temporal consistency via noise initialization, FIFO-Diffusion (kim2024fifo) feeds frames sequentially into a denoising window of training length, and Video-Infinity (tan2024video) exploits distributed computation to scale up video length. While effective for generating long videos, these methods are orthogonal to our length extrapolation strategy, which extends the intrinsic capacity of DiTs to longer sequences and can be readily integrated with them.
In addition to diffusion-based approaches to long video generation, alternative modeling paradigms such as autoregressive methods (wu2021godiva; yan2021videogpt; hong2022cogvideo; wu2022nuwa; kondratyuk2023videopoet; wu2024vila; sun2024generative; wang2024emu3) and diffusion forcing (chen2024diffusion; huang2025self; teng2025magi) are also capable of generating long videos. Although our method is designed for diffusion models, it may also offer insights into length extrapolation for these alternative paradigms.
@@ B.2 Details of the Multimodal Rotary Position Embedding
In this section, we provide the details of the Multimodal RoPE (M-RoPE) (wang2024qwen2) introduced in Sec. 2. Specifically, for a token at position $(t,h,w)$, the input vector ${\bm{x}}\in\mathbb{R}^{D}$ is divided into three subspaces of dimensions $d_{\mathcal{T}},d_{\mathcal{H}},d_{\mathcal{W}}$, respectively assigned to temporal, height, and width encodings. Each subspace is modulated by its own frequency series $\{\phi_{i}^{\mathcal{T}}\}_{i=0}^{d_{\mathcal{T}}-1},\{\phi_{i}^{\mathcal{H}}\}_{i=d_{\mathcal{T}}}^{d_{\mathcal{T}}+d_{\mathcal{H}}-1},\{\phi_{i}^{\mathcal{W}}\}_{i=d_{\mathcal{T}}+d_{\mathcal{H}}}^{D-1}$. Concretely, we define
@@ B.6 Remarks on Proposition 1
Model
Attention maps
Statistical row attention analysis
Hun.
Refer to caption Refer to caption

### v2

@@ 1 Introduction
Building upon the expressive power of diffusion transformers (DiTs) (Bao et al., 2023; Peebles and Xie, 2023), recent advances in text-to-video (T2V) generation (Bao et al., 2024; Zheng et al., 2024b; Brooks et al., 2024; Wan et al., 2025; Kong et al., 2024; Hong et al., 2022) have enabled models to synthesize high-fidelity videos. However, these models are typically trained on a fixed maximum sequence length (e.g., 5 seconds) (Wan et al., 2025; Kong et al., 2024; Hong et al., 2022) and struggle to generate videos beyond their training length, a task we term video length extrapolation, which is critical for practical applications.
To investigate the core challenges of this task, we conduct experiments on a range of models and identify two failure modes: (i) a model-specific periodic content repetition, where short clips loop indefinitely in certain models; and (ii) a universal quality degradation, manifested as blurred spatial details and frozen temporal dynamics across all models. Both failures become increasingly severe as the extrapolation length grows. Prior work, such as RIFLEx (Zhao et al., 2025), tackles repetition from the perspective of positional encodings, while overlooking quality degradation and therefore achieving limited extrapolation. We contend, however, that positional encodings play only an indirect role by perturbing queries and keys to influence attention. In contrast, attention itself—directly aggregating contextual information to generate outputs—offers a more fundamental view.
@@
To validate our approach, we conduct comprehensive evaluations on various T2V models (Kong et al., 2024; Yang et al., 2024; Wan et al., 2025) and extrapolation ratios, against a large family of baselines (Chen et al., 2023b; bloc97, 2023; Zhuo et al., 2024; Peng et al., 2023; Zhao et al., 2025). Experiments demonstrate that our method consistently surpasses all baselines in all settings by simultaneously addressing both failure modes. Notably, while prior methods collapse beyond $3\times$ extrapolation and yield static videos, ours maintains fluid motion, effectively extending the practical limit from $2\times$ to $4\times$. Remarkably, it improves Dynamic Degree and Imaging Quality by 233% and 40.5% over the previous best method at $4\times$ extrapolation. Beyond this, our method also generalizes seamlessly to downstream tasks such as various controllable video synthesis and editing.
HunyuanVideo Wan
Normal length Refer to caption Refer to caption
@@
3 $\times$ extra. Refer to caption Refer to caption
@@
Variable extra. Refer to caption
@@ 2 Preliminary
Attention mechanism with rotary position embedding. Modern video diffusion models are largely built on DiTs whose core is the attention mechanism (Vaswani et al., 2017; Li et al., 2025a). The input video is patched into $L$ tokens, each projected into queries, keys, and values. To encode the position information, DiTs mainly adopt Rotary Position Embedding (RoPE) (Su et al., 2024), which injects position into queries and keys through complex rotations. Concretely, for each query or key vector ${\bm{x}}\in\mathbb{R}^{D}$ at position $t$, RoPE maps it to $\mathbb{R}^{D}$ as
@@
For videos with temporal and spatial axes, Multimodal RoPE (M-RoPE) (Wang et al., 2024a) partitions the dimension $D=d_{\mathcal{T}}+d_{\mathcal{H}}+d_{\mathcal{W}}$ and encodes each subspace separately. Since we focus on temporal extrapolation, we consider only the temporal axis and denote $d_{\mathcal{T}}$ as $d$ for simplicity (see details in Appendix B.2).
Problem setting: video length extrapolation. Despite advances, DiT-based video generation models struggle to produce videos longer than their training duration. This task, known as video length extrapolation  (Zhao et al., 2025), aims to adapt a pre-trained model to generate high-quality videos of a sequence length $L^{\prime}$ that exceeds its training length $L$, with the extrapolation ratio defined as $s=L^{\prime}/L>1$. Notably, video length extrapolation targets the model’s intrinsic ability to generate longer sequences in a single forward generation, which is orthogonal to prior methods (Qiu et al., 2023; Wang et al., 2023; Kim et al., 2024; Wang et al., 2024c; Lu et al., 2024) that rely on inference-time modifications. See Appendix A for more related work.
@@ 3.1 Failure Modes of Video Length Extrapolation
In this section, we investigate the core challenges of video length extrapolation on a range of SOTA video diffusion transformers, including Wan (Wan et al., 2025), HunyuanVideo (Kong et al., 2024), and CogVideoX (Yang et al., 2024).
Qualitative results in Fig.2a and Fig.2b reveal two distinct failure modes. The first is a periodic content repetition, which occurs in certain models such as HunyuanVideo and CogVideoX. The second is a universal quality degradation, characterized by compromised spatial fidelity and temporal dynamics across all models. To further investigate their trends across extrapolation lengths, we perform a quantitative analysis on 10 prompts using metrics including Imaging Quality (Huang et al., 2024), Dynamic Degree (Huang et al., 2024), and Repetition Count. Fig. 2c confirms that both failures become more severe as the extrapolation factor increases.
@@ 3.2.1 The Cause of Content Repetition: Periodic Attention Patterns
Model Attention maps Statistical row-wise attention analysis
Hun. Refer to caption Refer to caption
@@
Wan Refer to caption Refer to caption
@@
Model Generated videos: baseline vs. intervention Attention maps: baseline vs. intervention
Hun. Refer to caption Refer to caption
@@
Wan Refer to caption Refer to caption
@@ 3.3 UltraViCo
Efficient CUDA implementation. UltraViCo requires modifying attention logits, but standard PyTorch attention is infeasible for long sequences. At a $3\times$ extrapolation ($\sim$200K tokens for HunyuanVideo), for instance, materializing a $200\text{K}\times 200\text{K}$ attention mask consumes over 8080GB of memory in bf16, causing an immediate out-of-memory error. To address this, we integrate UltraViCo into Triton-based FlashAttention (Dao et al., 2022) and SageAttention (Zhang et al., 2024b), where the online-softmax formulation avoids explicit mask construction. This yields scalable, memory-efficient computation, enabling UltraViCo on large video models.
@@ 4.1 Setup
We evaluate methods on three video diffusion models, including HunyuanVideo, Wan2.1-1.3B and CogVideoX-5B. Following RIFLEx, we use 100 prompts sampled from VBench (Huang et al., 2024). For quantitative evaluation, following RIFLEx, we adopt Imaging Quality (Quality), Dynamic Degree (Dynamics), and Overall Consistency (Overall) from VBench, along with the NoRepeat Score for models prone to content repetition. Notably, our NoRepeat Score is a variant of that in RIFLEx, tailored for multiple-repetition (see Appendix C.2 for details). Finally, we conduct a user study with 10 participants on 10 prompts, where users rank (User) the overall quality of videos across all methods. More details are provided in Appendix C.2.
@@ 4.2 Results
Performance comparison. We compare a wide range of length extrapolation baselines on three SOTA models (Kong et al., 2024; Yang et al., 2024; Wan et al., 2025) across various extrapolation ratios, including PE (Zhao et al., 2025), PI (Chen et al., 2023b), NTK (bloc97, 2023), TASR (Zhuo et al., 2024), YaRN (Peng et al., 2023), and RIFLEx. Tab. 1 reports $3\times$ and $4\times$ results on HunyuanVideo and Wan, while Fig. 6 shows qualitative samples on HunyuanVideo. Results for additional ratios and models are provided in the Appendix C.3.
@@
Connection with other long-video generation methods. UltraViCo aims to extend the effective training window of video diffusion transformers and is therefore orthogonal to existing long-video generation techniques such as FreeNoise (Qiu et al., 2023), FIFO-Diffusion (Kim et al., 2024), and sliding-window. As demonstrated in Table 2, enlarging the context window via UltraViCo consistently improves the long-term temporal consistency of these methods, without negatively affecting other performance. In Table 2, all methods follow the same evaluation setup ($6\times$ extrapolation for 30-second videos on Wan), where UltraViCo extends the base model’s training window by $3\times$.
Generalization to downstream tasks. Our method enhances the model’s inherent ability to handle longer sequences, making it naturally applicable to downstream tasks. As shown in Fig. 1, based on VACE (Jiang et al., 2025b), UltraViCo enables $3\times$ extrapolation in controllable generation and video editing. See Appendix C.4 for additional results.
@@ 5 Conclusion
Acknowledgements
This work was supported by Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (No. JYB2025XDXM101); National Natural Science Foundation of China (Nos. 62522609, 92470118); the Beijing Natural Science Foundation (No. L247030); and the fund for building world-class universities (disciplines) of Renmin University of China.
@@ Appendix A Related Work
Beyond autoregressive approaches Yan et al. (2025a; b), the recent advances in text-to-video generation have been primarily driven by diffusion models (Ho et al., 2020; Song et al., 2020; Ho et al., 2022; He et al., 2022; Zhao et al., 2022; 2023; Blattmann et al., 2023; Xing et al., 2023; Chen et al., 2023a; Zhao et al., 2024; Polyak et al., 2024; Zhou et al., 2024; Team, 2024a; Chen et al., 2024b; Ma et al., 2024a; b; 2025). With the development of diffusion transformers (DiTs) (Bao et al., 2023; Peebles and Xie, 2023), DiT-based text-to-video diffusion models have achieved remarkable performance, such as Sora (Brooks et al., 2024), Vidu (Bao et al., 2024), CogVideoX (Yang et al., 2024) and Open-Sora (Zheng et al., 2024a). Although achieving high quality, leading models are trained only on a fixed maximum sequence length, limiting long-term capacity. During video length extrapolation, they suffer from repetition or quality degradation, underscoring the need for length extrapolation.
@@
The goal of length extrapolation is to enable transformers to generate sequences longer than those seen during training in a single forward (Press et al., 2021). This is typically achieved by modifying positional encodings. For example, position interpolation (PI) (Chen et al., 2023b) improves performance by interpolating the frequencies in RoPE so that they remain within the training range even under extrapolation. NTK (bloc97, 2023), YaRN (Peng et al., 2023), and Time-aware Scaled RoPE (TASR) (Zhuo et al., 2024) combine interpolation with direct extrapolation, incorporating adjustments along the token dimension, denoising timesteps, and other factors to achieve better results. However, these methods perform poorly on image and video DiTs, often leading to content collapse or repetition. RIFLEx (Zhao et al., 2025) mitigates repetition by identifying and attenuating the intrinsic RoPE frequency, yet it still suffers from degraded visual quality. In contrast, our method effectively addresses both content repetition and quality degradation.
@@
There also exist many approaches to long video generation (Qiu et al., 2023; Wang et al., 2023; Henschel et al., 2025; Kim et al., 2024; Tan et al., 2024; Yin et al., 2025; Wang et al., 2024c; Cai et al., 2025; Li et al., 2025b; Lu et al., 2024; Tan et al., 2025; Jiang et al., 2025a; Gao et al., 2025; Gu et al., 2025), most of which intervene in the diffusion inference process. For instance, FreeNoise (Qiu et al., 2023) enhances temporal consistency via noise initialization, FIFO-Diffusion (Kim et al., 2024) feeds frames sequentially into a denoising window of training length, and Video-Infinity (Tan et al., 2024) exploits distributed computation to scale up video length. While effective for generating long videos, these methods are orthogonal to our length extrapolation strategy, which extends the intrinsic capacity of DiTs to longer sequences and can be readily integrated with them.
In addition to diffusion-based approaches to long video generation, alternative modeling paradigms such as autoregressive methods (Wu et al., 2021; Yan et al., 2021; Hong et al., 2022; Wu et al., 2022; Kondratyuk et al., 2023; Wu et al., 2024; Sun et al., 2024; Wang et al., 2024b) and diffusion forcing (Chen et al., 2024a; Huang et al., 2025; Teng et al., 2025) are also capable of generating long videos. Although our method is designed for diffusion models, it may also offer insights into length extrapolation for these alternative paradigms.
@@ B.2 Details of the Multimodal Rotary Position Embedding
In this section, we provide the details of the Multimodal RoPE (M-RoPE) (Wang et al., 2024a) introduced in Sec. 2. Specifically, for a token at position $(t,h,w)$, the input vector ${\bm{x}}\in\mathbb{R}^{D}$ is divided into three subspaces of dimensions $d_{\mathcal{T}},d_{\mathcal{H}},d_{\mathcal{W}}$, respectively assigned to temporal, height, and width encodings. Each subspace is modulated by its own frequency series $\{\phi_{i}^{\mathcal{T}}\}_{i=0}^{d_{\mathcal{T}}-1},\{\phi_{i}^{\mathcal{H}}\}_{i=d_{\mathcal{T}}}^{d_{\mathcal{T}}+d_{\mathcal{H}}-1},\{\phi_{i}^{\mathcal{W}}\}_{i=d_{\mathcal{T}}+d_{\mathcal{H}}}^{D-1}$. Concretely, we define
@@ B.6 Remarks on Proposition 1
Model Attention maps Statistical row attention analysis
Hun. Refer to caption Refer to caption